# Enhancing nuclear cross-section predictions with deep learning: the DINo algorithm

### v1

@@ Abstract
Accurate modeling of nuclear reaction cross-sections is crucial for applications such as hadron therapy, radiation protection, and nuclear reactor design. Despite continuous advancements in nuclear physics, significant discrepancies persist between experimental data and theoretical models such as TENDL (TALYS-based Evaluated Nuclear Data Library) [1], and ENDF/B (Evaluated Nuclear Data File) [2]. These deviations introduce uncertainties in Monte Carlo simulations widely used in nuclear physics and medical applications. In this work, DINo (Deep learning Intelligence for Nuclear reactiOns) is introduced as a deep learning-based algorithm designed to improve cross-section predictions by learning correlations between charge-changing and total cross-sections.
Trained on the TENDL-2021 dataset and validated against experimental data from the EXFOR database [3], DINo demonstrates a significant improvement in predictive accuracy over conventional nuclear models. The results show that DINo systematically achieves lower $\chi^{2}$ values compared to TENDL-2021 across multiple isotopes, particularly for proton-induced reactions on a 12C target, a key material in hadron therapy [4]. Specifically, for 11C production, DINo reduces the discrepancy with experimental data by $\sim 28\%$ compared to TENDL-2021. Additionally, DINo provides improved predictions for other relevant isotopes produced, such as 4He, 6Li, 9Be, and 10B, which play a crucial role in modeling nuclear fragmentation processes.
@@ 1 Introduction
Accurate modeling of nuclear reactions is essential for numerous applications, including hadron therapy, nuclear reactor design, and radiation protection. However, despite continuous advancements in nuclear physics, persistent discrepancies between experimental data and theoretical nuclear models remain a significant challenge, particularly in the context of particle therapy, where precise cross-section predictions directly impact treatment planning and dosimetry [7].
Traditionally, nuclear models such as TENDL (TALYS-based Evaluated Nuclear Data Library) [1] and ENDF/B (Evaluated Nuclear Data File) [2] rely on a combination of theoretical calculations and experimental benchmarking. However, these models often exhibit systematic deviations from measured cross-sections. Consequently, Monte Carlo (MC) simulations—widely used for predicting nuclear interactions in medical and nuclear physics applications—face inherent accuracy limitations due to uncertainties in nuclear data.
A conventional approach to mitigating these inaccuracies involves integrating experimental cross-section measurements to refine Monte Carlo simulations [4]. While this method enhances predictive reliability, it remains dependent on the availability of high-quality experimental data, which is often limited for certain isotopes and energy ranges. An alternative strategy is to directly enhance the predictive capabilities of nuclear models by incorporating machine learning techniques capable of capturing complex patterns within nuclear data. Notably, [Henning] have investigated similar approaches for inelastic neutron scattering cross-sections. Part of the results and analyzes presented here were initially developed in the PhD thesis [8].
In this work, DINo (Deep learning Intelligence for Nuclear reactiOns) is introduced as a deep learning-based algorithm designed to address these challenges. DINo leverages neural networks to learn the underlying correlations between partial nuclear reaction cross-sections and total cross-sections, thereby improving the accuracy of nuclear reaction predictions beyond what is achievable with conventional nuclear models. Specifically, DINo is trained using TENDL-2021 nuclear data, with the objective of predicting as a starting point of the total cross-sections for proton-induced nuclear reactions on a 12C target. The choice of protons is motivated by their prominent role in hadron therapy, where accurate cross-section estimations are crucial for dose calculations and treatment optimization.
@@ 2 Material and methods
2.1 TENDL-2021 nuclear model
TENDL (TALYS-based Evaluated Nuclear Data Library) [1] is a nuclear data library developed by the Nuclear Reaction Data Centre (NRDC) at the Institute for Reference Materials and Measurements (IRMM), a part of the Joint Research Centre (JRC) of the European Commission. While originally based on the TALYS theoretical nuclear reaction code, the version used in this study incorporates evaluations from the ENDF/B (Evaluated Nuclear Data File) library [2], first distributed with ENDF/B-VII.0 and still available in ENDF/B-VIII.0.
Nuclear data libraries like ENDF/B and TENDL provide essential information on nuclear reactions, including cross-sections, decay data, and other fundamental nuclear properties. These are critical for applications such as nuclear reactor design, nuclear medicine, and nuclear security. Each new release of TENDL integrates refinements in theoretical modeling, updated experimental data, and corrections based on benchmarking against experimental results and user feedback.
In this study, protons were selected due to their extensive coverage in nuclear data libraries. To ensure a uniform energy step distribution and standardize the ENDF6 format 10.2172/2007538, an algorithm developed by J. Collin (private communication) was applied.
2.2 EXFOR experimental data
@@
The database compiles measured nuclear cross-sections, energy spectra, and angular distributions obtained from worldwide experiments. It standardizes data formats for easy integration into nuclear models. EXFOR gathers data from a wide range of experimental facilities, including accelerators, reactors, and other nuclear research facilities. Data contributors can include universities, research institutes, and national laboratories. The data are publicly accessible and can be accessed through various channels, including the EXFOR database website (https://www-nds.iaea.org/exfor/), which provides search and retrieval functionalities. Users can search for specific reactions, isotopes, energy ranges, and other parameters to retrieve relevant experimental data.
Since proton-induced reactions are well-documented in EXFOR, this dataset was chosen to evaluate and validate our deep learning model, providing a direct comparison between experimental and theoretical nuclear cross-sections.
2.3 Proton beam interaction with 12C target
As mentioned in the previous section, the focus will be on proton interaction, and more especially on a proton beam on a 12C target. Figure 1 shows the total and partial cross-sections for incident protons on a 12C target from the TENDL-2021 nuclear model alongside the partial ones. Indeed, this choice is particularly relevant in the context of nuclear reactions in hadrontherapy.
@@ 2.3 Proton beam interaction with 12C target
Firstly, 12C is often referred to as a ”tissue-equivalent” material, meaning it has similar nuclear properties to biological tissues. This characteristic makes it an ideal target for simulating interactions that occur in human tissues during therapeutic procedures, such as in hadron therapy.
Secondly, the interaction of a proton beam with a 12C target can produce secondary particles such as protons, neutrons, and fragments of carbon nuclei coming from the target. These secondary particles are crucial for understanding the dosimetry and biological effects of the radiation delivered to the patient. This choice of using a proton projectile on a 12C target is justified because, in modern nuclear physics experiments that measure cross-sections, such as the FOOT experiment [4], direct measurement of carbon ion fragmentation in tissue is complex and challenging. This approach leverages the symmetry of nuclear interactions: the fragments produced when a proton hits a 12C nucleus are similar to those produced when a 12C ion hits a hydrogen nucleus (which is a major component of human tissue). This makes proton-induced reactions an indirect yet effective method for studying nuclear fragmentation in medical applications.
2.4 Deep learning algorithms
@@ 2.5 DINo structure
The primary goal of DINo is to predict total cross-sections for initial energies that are not covered by current experimental data. This is crucial in hadrontherapy where accurate treatment plan predictions rely on understanding nuclear reactions and their implications. Figure 2 depicts the global structure of the DINo algorithm. The choices made for DINo model architecture are explained and justified in Annexe of the PhD thesis LABEL:Gesson.
@@
After undergoing 20 epochs of training, the DINo algorithm is ready for application to experimental data. This transition from theoretical modeling to experimental prediction is a critical step in validating the algorithm’s effectiveness and real-world applicability.
During the testing phase on experimental data, the algorithm receives actual values for charge-changing cross-sections along with their associated initial energies. Leveraging the knowledge gained during the learning phase, DINo can then predict the corresponding experimental values for total cross-sections at these specific initial energies. This predictive capability is essential in scenarios where direct experimental measurements are limited or impractical, allowing to extrapolate crucial information for various nuclear processes in simulation for example.
2.6 Comparison of DINo predictions with the TENDL Model
@@ 3.1 Detailed Study of 11C Product Cross-Section
Figure 3 presents the total cross-section and 11C product cross-section for incident protons on a 12C target, as a function of incident energy. The TENDL-2021 model predictions are plotted alongside experimental data to assess their agreement.
The $\chi^{2}$ evaluation of the TENDL-2021 model’s total cross-section predictions relative to experimental data yields a value of:
@@
which indicates a moderate but quite notable deviation from the experimental dataset.
@@
Figure 4: Algorithm test on total cross-section for incident protons on a 12C target as a function of incident energy: comparison of TENDL model results with DINo predictions.
@@ 3.2 Extension to other isotopic products
Table 1 summarizes the $\chi^{2}$ values comparing DINo’s predictions to TENDL-2021 across these isotopes.
@@
Across all isotopes studied, DINo systematically achieves lower $\chi^{2}$ values, confirming that it predicts outcomes more accurately than TENDL-2021.
@@ 4 Discussion
A key advantage of DINo over traditional nuclear models like TENDL-2021 is its computational efficience. Once trained, the deep learning model produces cross-section predictions in the order of microseconds, whereas TENDL relies on extensive simulations, requiring significantly more computational time. This real-time prediction capability of DINo makes it particularly advantageous for applications requiring fast nuclear data processing like hadron therapy treatment planning, or accelerator-based experiments.
Despite its advantages, DINo’s performance remains highly dependent on the quality of training data. The accuracy of predictions deteriorates in energy regions with limited experimental data (e.g., below 10 MeV and above 100 MeV), where the models perform less well and where fewer data are available for statistical evaluation of the predictions. This observation suggests that:
@@
The ability to extrapolate nuclear cross-sections across isotopes suggests that deep learning models like DINo could eventually be used as broad-spectrum nuclear data predictors, offering a complementary tool to traditional evaluated nuclear data libraries.

### v2

@@ Abstract
Accurate modeling of nuclear reaction cross-sections is crucial for applications such as hadron therapy, radiation protection, and nuclear reactor design. Despite continuous advancements in nuclear physics, significant discrepancies persist between experimental data and theoretical models such as TENDL (TALYS-based Evaluated Nuclear Data Library), and ENDF/B (Evaluated Nuclear Data File). These deviations introduce uncertainties in Monte Carlo simulations widely used in nuclear physics and medical applications. In this work, DINo (Deep learning Intelligence for Nuclear reactiOns) is introduced as a deep learning-based algorithm designed to improve cross-section predictions by learning correlations between charge-changing and total cross-sections.
Trained on the TENDL-2021 dataset and validated against experimental data from the EXFOR database, DINo demonstrates a significant improvement in predictive accuracy over conventional nuclear models. The results show that DINo systematically achieves lower $\chi^{2}$ values compared to TENDL-2021 across multiple isotopes, particularly for proton-induced reactions on a 12C target, a key material in hadron therapy. Specifically, for 11C production, DINo reduces the discrepancy with experimental data by $\sim 28\%$ compared to TENDL-2021. Additionally, DINo provides improved predictions for other relevant isotopes produced, such as 4He, 6Li, 9Be, and 10B, which play a crucial role in modeling nuclear fragmentation processes.
@@ 1 Introduction
Accurate modeling of nuclear reactions is essential for numerous applications, including hadron therapy, nuclear reactor design, and radiation protection. However, despite continuous advancements in nuclear physics, persistent discrepancies between experimental data and theoretical nuclear models remain a significant challenge, particularly in particle therapy, where precise cross-section predictions directly impact treatment planning and dosimetry [1].
Traditionally, nuclear models such as TENDL (TALYS-based Evaluated Nuclear Data Library) [2] and ENDF/B (Evaluated Nuclear Data File) [3] rely on a combination of theoretical calculations and experimental benchmark. However, these models often exhibit systematic deviations from measured cross-sections. Consequently, Monte Carlo (MC) simulations—widely used for predicting nuclear interactions in medical and nuclear physics applications—inherit of this biaised data.
A conventional approach to mitigating these inaccuracies involves integrating experimental cross-section measurements to refine Monte Carlo simulations [4]. While this method enhances predictive reliability, it remains dependent on the availability of high-quality experimental data, which is often limited for certain isotopes and energy ranges. An alternative strategy is to directly enhance the predictive capabilities of nuclear models by incorporating machine learning techniques capable of capturing complex patterns within nuclear data. Such attempts have been investigate in a similar approach for inelastic neutron scattering cross-sections [Henning].
In this work, initially developed in the PhD thesis [5], DINo (Deep learning Intelligence for Nuclear reactiOns) is introduced as a deep learning-based algorithm designed to address these challenges. DINo leverages neural networks to learn the underlying correlations between partial nuclear reaction cross-sections and total cross-sections, thereby improving the accuracy of nuclear reaction predictions beyond what is achievable with conventional nuclear models. Specifically, DINo is trained using TENDL-2021 nuclear data, to predict as a starting point of the total cross-sections for proton-induced nuclear reactions on a 12C target. The choice of protons is motivated by their prominent role in hadron therapy, where accurate cross-section estimations are crucial for dose calculations and treatment optimization.
@@ 2 Material and methods
2.1 TENDL-2021 nuclear model and EXFOR experimental data
TENDL (TALYS-based Evaluated Nuclear Data Library) [2] is a nuclear data library produced at NRG Petten. Since 2015, TENDL is mainly developed at PSI and the IAEA (Nuclear Data Section). While originally based on the TALYS theoretical nuclear reaction code, the main version distributed incorporates data from the other evaluator ENDF/B (Evaluated Nuclear Data File) library [3], first distributed with ENDF/B-VII.0 and still available in ENDF/B-VIII.0. This remains the case for the reaction of interest.
Nuclear data libraries like ENDF/B and TENDL provide essential information on nuclear reactions, including cross-sections, decay data, and other fundamental nuclear properties. Each new release of TENDL integrates refinements in theoretical modeling, updated experimental data, and corrections based on benchmarking against experimental results and user feedback.
In this study, protons were selected due to their extensive coverage in nuclear data libraries. To ensure a uniform energy step distribution and standardize the ENDF6 [6], an algorithm developed by J. Collin (private communication) was applied.
@@ 2.1 TENDL-2021 nuclear model and EXFOR experimental data
The database compiles measured nuclear cross-sections, energy spectra, and angular distributions obtained from worldwide experiments. It standardizes data formats for easy integration into nuclear models. EXFOR gathers data from a wide range of experimental facilities, including accelerators, reactors, and other nuclear research facilities. Data contributors can include universities, research institutes, and national laboratories. The data are publicly accessible and can be accessed through various channels, including the EXFOR database website, which provides search and retrieval functionalities. Users can search for specific reactions, isotopes, energy ranges, and other parameters to retrieve relevant experimental data.
Since proton-induced reactions are well-documented in EXFOR, this dataset was chosen to evaluate and validate our deep learning model, and more especially a proton beam on a 12C target, providing a direct comparison between experimental and theoretical nuclear cross-sections. Figure 1 shows the total and partial cross-sections for incident protons on a 12C target from the TENDL-2021 nuclear model.
@@
Firstly, 12C is often referred to as main constituent of ”tissue-equivalent” material, meaning it has similar nuclear properties to biological tissues. This characteristic makes it an ideal target for simulating interactions that occur in human tissues during therapeutic procedures, such as in hadron therapy.
Secondly, the interaction of a proton beam with a 12C target can produce secondary particles such as protons, neutrons, and fragments of carbon nuclei coming from the target. These secondary particles are crucial for understanding the dosimetry and biological effects of the radiation delivered to the patient. This choice of using a proton projectile on a 12C target is justified because, in modern nuclear physics experiments that measure cross-sections, such as the FOOT experiment [4], direct measurement of carbon ion fragmentation in tissue is complex and challenging. This approach, inverse cinematic of carbon fragmentation, leverages the symmetry of nuclear interactions: the fragments produced when a proton hits a 12C nucleus are similar to those produced when a 12C ion hits a hydrogen nucleus (which is a major component of human tissue). This makes proton-induced reactions an indirect yet effective method for studying nuclear fragmentation in medical applications.
2.2 DINo deep learning algorithm
@@
The primary goal of DINo is to predict total cross-sections for initial energies that are not covered by current experimental data. This is crucial in hadrontherapy where accurate treatment plan predictions rely on understanding nuclear reactions and their implications. Figure 2 depicts the global structure of the DINo algorithm. The DINo model architecture choices are explained and justified in Annexe of the PhD thesis [5].
@@
During the second testing phase, on experimental data, the algorithm receives actual experimental values for charge-changing cross-sections along with their associated initial energies. Leveraging the knowledge gained during the learning phase, DINo can then predict the corresponding experimental values for total cross-sections at these specific initial energies. This predictive capability is essential in scenarios where direct experimental measurements are limited or impractical, allowing extrapolation of crucial information for various nuclear processes in simulation for example.
@@ 3.1 Detailed Study of 11C Product Cross-Section
Figure 3 presents the total cross-section and 11C product cross-section for incident protons on a 12C target, as a function of incident energy. The TENDL-2021 model predictions are plotted alongside experimental data to assess their agreement. The $\chi^{2}$ evaluation of the TENDL-2021 model’s total cross-section predictions relative to experimental data yields a value of:
@@
which indicates a quite notable deviation from the experimental dataset.
@@
Figure 4: Algorithm test on the total cross-section for incident protons on a 12C target as a function of incident energy: comparison of TENDL model results with DINo predictions.
@@ 3.2 Extension to other isotopic products
Table 1 summarizes the $\chi^{2}$ values comparing DINo’s predictions to TENDL-2021 across these isotopes. Across all isotopes studied, DINo systematically achieves lower $\chi^{2}$ values, confirming that it predicts outcomes more accurately than TENDL-2021.
@@ 4 Discussion
A key advantage of DINo over traditional nuclear models like TENDL-2021 is its computational efficiency. Once trained, the deep learning model produces cross-section predictions in the order of microseconds, whereas TENDL relies on extensive simulations, requiring significantly more computational time. This real-time prediction capability of DINo makes it particularly advantageous for applications requiring fast nuclear data processing like hadron therapy treatment planning, or accelerator-based experiments.
Despite its advantages, DINo’s performance remains highly dependent on the quality of training data. The accuracy of predictions deteriorates in energy regions with limited experimental data (e.g., below 10 MeV and above 100 MeV), where the models perform less well and where less data are available for statistical evaluation of the predictions. This observation suggests that:
@@
The ability to extrapolate nuclear cross-sections across isotopes suggests that deep learning models like DINo could eventually be used as broad-spectrum nuclear data predictors, offering a complementary tool to traditionally evaluated nuclear data libraries.
@@ 5 Conclusion
Data Availability Statement: No Data associated in the manuscript